OpenStudy (anonymous):

please help me with this: ( linear differential equations) ydx-4(x+y^6)dy=0

OpenStudy (freckles):

is the differential equation exact?

OpenStudy (anonymous):

I have to do it for the method of linear differential equations

OpenStudy (freckles):

well what does that exactly mean... Have you not talked about exact differential equations?

OpenStudy (anonymous):

no, that is a diferent method, linear differential equations: dy/dx+p(x)y=f(x)

OpenStudy (freckles):

so can you put it in that form then?

OpenStudy (freckles):

I don't think you can but if you can I will help you further.

OpenStudy (freckles):

\[y-4(x+y^6)y'=0 \\ -4(x+y^6)y'+y=0 \\ y'+\frac{-1}{4(x+y^6)}y=0\] see how that isn't in that form?

OpenStudy (freckles):

\[y'+p(x,y)y=f(x)\]

OpenStudy (anonymous):

it sould be just P(X) not P(X,Y)

OpenStudy (freckles):

well then can you show me how you get p(x) there then

OpenStudy (freckles):

because I don't see how

OpenStudy (anonymous):

that is the problem that I have haha

OpenStudy (freckles):

You can't put the problem in y'+p(x)y=f(x) form

OpenStudy (freckles):

the thing next to the y is p(x,y) not p(x)

OpenStudy (freckles):

there has to be some other way you have learned besides putting it in that form and basing the integrating factor off that form

OpenStudy (freckles):

also you called this a linear diff equation it is a nonlinear